OpenStudy (anonymous):

A 1300 kg sports car accelerates from rest to 95 km/h in 7.4 s. What is the average power delivered by the engine?

OpenStudy (anonymous):

Power = work done / time work done = Change in K.E hope this helps/......

OpenStudy (ghazi):

\[F=m*a\] to find a use \[a=\frac{ V }{ t }=\frac{ 26.38 }{ 7.4 }=3.566m/s^2\] now F= 1300*3.566=4635.88N now find the distance covered in 7.4 seconds \[D=ut+\frac{ 1 }{ 2 }at^2\] u=0 , D= 0.5*3.566*(7.4)^2=97.637m now work done = F*D=4635.88*97.637=452633.786 j finally power = w/t= 452633.768/7.4=61166.727 Watts looks like its bugati veyron lol

OpenStudy (ghazi):

damn :( lol

OpenStudy (anonymous):

work done = Change in K.E =1/2 m v^2 - 0

OpenStudy (ghazi):

correct